# A Multimodal LLM Approach for Visual Question Answering on Multiparametric 3D Brain MRI

### v2

@@ 3.2.1 Synthetic VQA protocol
We use the conform function from the Nibabel python library (Abraham et al., 2014) to put the BraTS volumes and the LPBA40 atlas (Shattuck et al., 2008) in the RAS space. We then overlay the atlas on each mask volume and, based on their intersection, extract the following brain regions: “frontal”, “parietal”, “occipital”, “temporal”, “limbic”, “insula”, “subcortical”, “cerebellum”, and “brainstem”.
@@
We identify all disconnected islands, noting the largest as the “core,” and compute what proportion of the total mask volume it occupies. If there is only one island, the pattern is “single lesion”; if multiple islands are present but the core retains $\geq$70% of the volume, it is described as “core with satellite lesions”; otherwise, when no dominant island exists, the distribution is marked “scattered lesions.”
@@
After computing the previous quantities for each label mask, we create a dataset that simulates the natural variability of human input. First, we consider all combinations of the four major tasks to create multi-task question-answer pairs. After we have the 15 question-answer pair types, we use ChatGPT-4o to generate approximately 3000 perturbations of each question-answer pair (without affecting the label and answer term) that emulates the language a clinician would use. We also add question-answer pairs with partially out-of-scope and completely out-of-scope tasks to improve the model’s self-awareness of its capabilities. Thus, for each label and mpMRI in each dataset, we sample four multitask question-answer pairs without replacement such that each major task is addressed in at least one question-answer pair, one partially out-of-scope question-answer pair, and one completely out-of-scope question-answer pair. Examples of generated question-answer pairs can be seen in the Appendix in Table 5. The answers are used as supervision for next-token prediction for the multimodal LLM.
3.2.2 Multi-task heads
@@
$\mathcal{L}=\mathcal{L}_{\textrm{Next-token}}+\mathcal{L}_{\textrm{Volume}}+\mathcal{L}_{\textrm{Region}}+\mathcal{L}_{\textrm{Shape}}+\mathcal{L}_{\textrm{Spread}}+\mathcal{L}_{\textrm{Out-of-scope}}$ (2)
4 Experiments
4.1 Datasets details
@@
For our synthetic VQA protocol, we leverage the Brain Tumor Segmentation (BraTS) challenge (LaBella et al., 2024), which provides a standardized benchmarking environment for automated brain tumor segmentation. All datasets comprise of co-registered multiparametric MRI scans (T1, T1Gd, T2, FLAIR) at 1mm3 resolution, skull-stripped and manually annotated by experts. To enable fair comparison and manage GPU memory, all BraTS sequences were resampled to $32\times 256\times 256$. This allowed for compatibility with baseline methods, such as M3D (Bai et al., 2024) and Med3DVLM (Xin et al., 2025). We consider three challenges in BraTS: GLI, MET and GoAT. The challenges are collected from over ten institutions and encompass diverse pathological contexts and imaging protocols.
@@
GLI (Adult Glioma Post Treatment) focuses on post-treatment diffuse glioma segmentation and consists of multi-institutional routine post-treatment clinically-acquired multiparametric mpMRI scans of glioma. The task requires the delineation of enhancing tumor (ET), non-enhancing tumor core (NETC), surrounding FLAIR hyperintensity (SNFH), and resection cavity (RC) (de Verdier et al., 2024).
@@
MET (Brain Metastases) contains a retrospective compilation of treatment-naive brain metastases mpMRI scans obtained from various institutions under standard clinical conditions. The challenge addresses the segmentation of small metastatic lesions using a 3-label system (NETC, SNFH, ET) and demonstrates variable tumor component distribution across cases (Moawad et al., 2024).
@@
GoAT (Generalizability Across Tumors) assesses algorithmic generalizability across different tumor types (i.e., different number of lesions per scan, lesion sizes, and locations in the brain), institutions (i.e., different MRI scanners, acquisition protocols), and demographics (i.e., different age, sex, etc.). The challenge uses consistent labels (necrosis, edema/invaded tissue, and enhancing tumor) despite varying tumor morphology to evaluate algorithm adaptability to new disease types with limited training data (de Verdier et al., 2024; Moawad et al., 2024; LaBella et al., 2023; Kazerooni et al., 2024; Adewole et al., 2023).
@@
To generate the train, validation, and test sets, we randomly sample 80%, 10%, and 10% from the imaging studies. For GLI we generated 31,104, 4,176, and 3,624 question-answer pairs for the train, validation, and test sets based on 1,621 mpMRIs. For MET, we generated 9,090, 1,368, and 1,260 question-answer pairs for the train, validation, and test sets, based on 651 mpMRIs. For GoAT, we generated 19,440, 2,430, and 2,448 question-answer pairs for the train, validation, and test sets, based on 1351 mpMRIs.
Clinical validation We collaborated with a radiologist who annotated 10 multiparametric MRIs (mpMRIs) from the BraTS-GLI test set, with questions spanning four tasks and four labels per case, yielding a total of 160 questions. We evaluated the inter-annotator agreement between the clinician and our synthetic annotations using the Kappa score, which was 50.4, indicating moderate agreement. To contextualize the value, inter-annotator segmentation agreement for substructures in post-operative glioma typically ranges between 30.0–40.0 Jaccard score (Visser et al., 2019), indicating the difficulty of this task. While no directly comparable study exists for glioma terminology, a relevant meta-analysis of expert inter-annotator agreement on breast cancer terminology reported a median mean Kappa score of 43.8 (Antonio & Crespi, 2010). Furthermore, to assess the quality of our synthetic questions, the same radiologist evaluated the clarity of 160 synthetic questions from the same 10 mpMRIs using a binary scoring system (1 = valid, 0 = invalid). The synthetic questions achieved a 92.2% validity rate, indicating high acceptability.
More statistics about the synthetic datasets can be seen in Table 1, and more details can be found in Appendix ABC.
Figure 3: Heatmap for correlation between high-level expert weight vectors for standard prompts in the GLI dataset. NETC = non-enhancing tumor core, ET = enhancing tissue, SNFH = surrounding FLAIR hyperintensity, RC = resection cavity.
4.2 Experimental settings
@@
We compare our approach to several baseline models, including LLaVA-Med (Li et al., 2023a), M3D (Bai et al., 2024), Med3DVLM (Xin et al., 2025), and RadFM (Wu et al., 2023) [^1] We planned to evaluate Merlin (Blankemeier et al., 2024), but the report-generation model weights were not publicly available at the time of submission.. To process the multiple 3D MRI image modalities, we use a multi-image approach, in which we concatenate the image tokens generated from each MRI image modality from a shared projection layer and vision encoder (Wu et al., 2023). Because LLaVA-Med is not implemented with a 3D vision encoder, to ensure a fair comparison, we test it with our model’s vision encoder (Bai et al., 2024). Similar to our method, the vision encoder is frozen and only the projection layer and LLM are trainable. To provide a comparison to our model’s multi-task heads, which are trained end-to-end with the rest of our framework, we independently train a new multi-task head. We use a Phi3 language model with multi-task heads to predict the multi-task outputs given the prompt and text generation. The model is trained on our train dataset and had 99.8% accuracy on the validation set. Other hyperparameter settings mirror our method as closely as possible to ensure a fair comparison.
@@
All our experiments were mainly conducted using a single NVIDIA A100 GPU on an internal cluster. Training our model on the GLI dataset took roughly 8 hours.
4.3 Results
Table 2: Comparison of task performance for all models on all datasets with accuracy metric with standard deviation.
@@ 4.3 Results
Table 3: Ablation study on the MoE architecture on the GLI validation set with accuracy metric.
[…]
rowsep=0.95pt Modality-level MoE Token-level MoE Prompt-based MoE weights Task Mean 68.7 69.1 69.9 70.4
Table 4: Model comparison with multi-task loss on the GLI validation set with accuracy metric.
[…]
rowsep=0.95pt, colsep=3.5pt Method Task Mean mpLLM without multi-task loss 67.4 mpLLM with multi-task loss 70.4
All model results across the evaluated datasets are presented in Table 2. Our model consistently achieves strong performance across all task categories and datasets, outperforming the second-best model by an average margin of 5.3%. Furthermore, it ranks first in nearly all sub-categories and datasets, highlighting both its broad capabilities and strong generalizability. We also validated model performance on the clinically annotated subset and graded the top two models’ responses for sufficiency, which can be seen in Table 10 and Table 11 respectively in the Appendix, and note that our model outperforms the baselines. Examining the memory usage, our model only required approximately 20 GB of GPU memory during training and inference – significantly less than M3D, LLaVA-Med, and Med3DVLM, all of which exceed 40 GB – suggesting the computational benefits of a fused vision token representation. In our experiments, we also noticed that the top three models had above a 99.8% accuracy on out-of-scope task identification, which suggests our dataset was effective at hallucination mitigation.
4.4 Ablation studies
The ablation study on the MoE architecture is in Table 3. Image modality-level and token-level high-level MoE experts perform better than the single projection layer baseline approach. A weighted combination of the different high-level experts conditioned on the text prompt performs the best, which is what our hierarchical MoE architecture utilizes.
To qualitatively evaluate our architecture, we construct all 60 template task prompts from our GLI dataset (four labels $\times$ 15 task combinations = 60 template prompts) and input them into our model’s high-level router to generate high-level expert weight vectors. We then calculate the correlation between these weight vectors and generate a heatmap, which is in Figure 3. There’s high correlation between expert weight vectors within the same label, suggesting similar image features are extracted. For labels that are closer anatomically, such as non-enhancing tumor core and enhancing tissue, there is also relatively high correlation between the expert weight vectors. This is reasonable because of their close proximity anatomically, which suggests similar extracted image features. For labels like resection cavity and surrounding FLAIR hyperintensity that are more diverse anatomically from the other labels, there’s much lower correlation, which again is sensible.
In Table 4, we see a comparison of our model performance trained with our multi-task loss versus the next-token prediction baseline loss. Clearly, there is a significant performance improvement with our multi-task loss.
5 Conclusion
We present mpLLM, a multimodal LLM with prompt-conditioned hierarchical MoE that routes across modality- and token-level projection experts for mpMRI VQA, enabling efficient end-to-end fine-tuning without paired image–report pretraining. With a clinician-validated synthetic VQA pipeline derived from segmentation annotations, mpLLM improves over strong medical VLM baselines by an average of +5.3% while using <<50% GPU memory. Ablations highlight the modality/token experts, prompt-conditioned routing, and an integrated multi-task head. Future work includes open-ended VQA/report generation, broader multi-reader validation, and fairness analyses.
6 Ethics statement
This work uses publicly available, fully de-identified BraTS datasets, minimizing risks to patient privacy and data security. Our synthetic VQA questions are generated from segmentation annotations, and both the generated questions and model outputs underwent clinician review to mitigate typical risks of synthetic supervision. Nonetheless, fairness and bias remain open concerns: synthetic prompts and limited demographic metadata can yield models that underperform for underrepresented groups or clinical scenarios. The model is intended for research only and must not be used for autonomous clinical decision-making; it is designed to abstain on out-of-scope queries, and any deployment would require prospective, multi-site validation under qualified clinical oversight. In future work, we will expand evaluations to demographically diverse cohorts where available, document dataset composition and known limitations, and incorporate explicit fairness analyses and bias-mitigation strategies alongside robustness and calibration assessments.
7 Reproducibility statement
We use a publicly available dataset and detail the full data-generation pipeline in Sections 3.2.1 and 4.1, with additional information in Appendix A. We report all experimental settings and computational resources in Section 4.2.
Appendix A Additional information regarding dataset
@@
In the following, we will describe the formulas used to derive the shape and spread descriptors for our synthetic VQA protocol. Let $M\subset\mathbb{Z}^{3}$ be a binary mask of foreground voxels sampled with spacing $\mathbf{s}=(s_{x},s_{y},s_{z})\,[\mathrm{mm}]$ (typically $s_{x}=s_{y}=s_{z}=1$). Write $\Delta V=s_{x}s_{y}s_{z}$ for the physical volume of one voxel and $\lvert M\rvert$ for the number of foreground voxels.
@@
We decompose $M$ into 26-connected components $M_{1},\dots,M_{N_{\!c}}$ (scipy ‘ndimage.label‘ with a unit “ball” structuring element) and record $N_{\!c}$.
@@
$\text{shape}=\begin{cases}\text{``focus''}&V_{\text{tot}}<0.1\;\text{cm}^{3}\quad(\;V_{\text{tot}}\times 10^{-3}<0.1\;)\\[2.0pt] \text{``round''}&\Phi\geq 0.85\;\land\;E\leq 1.3,\\[2.0pt] \text{``oval''}&0.60\leq\Phi<0.85\;\land\;1.3<E\leq 2.5,\\[2.0pt] \text{``elongated''}&E>2.5,\\[2.0pt] \text{``irregular''}&\text{otherwise.}\end{cases}$
The thresholds were set empirically on a development set of annotated masks and match clinicians’ qualitative intuition of near-spherical, mildly flattened, and strongly stretched geometries. All computations are implemented in Python using scipy, scikit-image, numpy, and ndimage as shown in the listing above.
@@
We use ChatGPT to generate question augmentations of our multitask dataset. For generating question augmentations for the standard multi-task prompts, we first provide this prompt “Please produce hundred alternative wordings that a clinician may use for the following question and answer. Please include everything surrounded by curly braces {} as they are because they are placeholders. Please generate the reworded question starting with “Q:” and reworded answer starting with “A:” and separate each generated question-answer pair with a newline. Please do not produce any additional text.” and append this to each of the multitask prompts below. We produce 40 repetitions with a temperature of $1.0$, top p of $1$, and model “gpt-4o-mini-2024-07-18”.
  1. 1.
@@
For generating question augmentations for the partially out-of-scope multi-task prompts, we first provide this prompt “Please produce hundred alternative wordings that a clinician may use for the following question and answer and incorporate an additional clinical task or tasks which the model cannot solve in the reworded question. These can be before, after, or interspersed between the other tasks (please make sure to vary the order and number of out-of-scope tasks). Do not mention that the model cannot answer these in the question; however, indicate that the model cannot answer that part of the question in the reworded answer (potentially using different phrasings). The model can describe the volume, brain region, shape, and spread of {label} which is the region of interest. Please include everything surrounded by curly braces {} as they are because they are placeholders. Please generate the reworded question starting with “Q:” and reworded answer starting with “A:” and do not produce any additional text.” and append this to each of the multitask prompts above. We produce 10 repetitions with a temperature of $1.0$, top p of $1$, and model “gpt-4o-2024-08-06”.
@@
For generating question augmentations for completely out-of-scope prompts, we first provide this prompt “Please produce a hundred questions (with one or more tasks) that a clinician may ask that the model does not have information to answer. The model can describe the volume, brain region, shape, and spread of {label} which is the region of interest. Please include {label} in the question but do not include anything else with curly braces. In the answer, please indicate the model cannot answer the question (potentially using different phrasings). Please generate the question starting with “Q:” and answer starting with “A:” and do not produce any additional text.” We produce 10 repetitions with a temperature of $1.0$, top p of $1$, and model “gpt-4o-mini-2024-07-18”.

### v3

@@ 3.2.1. Synthetic VQA protocol
For the BraTS GLI volumes, we use the Nibabel python library (Abraham et al., 2014) to register the volumes to the AAL atlas (version SPM12) (Rolls et al., 2020) and extract the following brain regions: “frontal”, “parietal”, “occipital”, “temporal”, “limbic”, “insula”, “subcortical”, and “cerebellum”. For the BraTS MET and GoAT volumes, we register the volumes to the LPBA40 atlas (in SRI24 space) (Shattuck et al., 2008) and extract the following brain regions: “frontal”, “parietal”, “occipital”, “temporal”, “limbic”, “insula”, “subcortical”, “cerebellum”, and “brainstem”. The percent coverages of the segmentation masks with the atlases are 67.3%, 70.9%, and 57.7% for the GLI, MET, and GoAT datasets respectively.
@@
We identify all disconnected islands, noting the largest as the “core,” and compute what proportion of the total mask volume it occupies. If there is only one island, the pattern is “single lesion”; if multiple islands are present but the core retains $\geq$85% of the volume, it is described as “core with satellite lesions”; otherwise, when no dominant island exists, the distribution is marked “scattered lesions.”
@@
After computing the previous quantities for each label mask, we create a dataset that simulates the natural variability of human input. First, we consider all combinations of the four major tasks to create multi-task question-answer pairs. After we have the 15 question-answer pair types, we use ChatGPT-4o to generate approximately 3000 perturbations of each question-answer pair (without affecting the label and answer term) that emulates the language a radiologist would use. We also add question-answer pairs with partially out-of-scope and completely out-of-scope tasks to improve the model’s self-awareness of its capabilities. Thus, for each label and mpMRI in each dataset, we sample four multitask question-answer pairs without replacement such that each major task is addressed in at least one question-answer pair, one partially out-of-scope question-answer pair, and one completely out-of-scope question-answer pair. Examples of generated question-answer pairs can be seen in the Appendix in Table 11. The answers are used as supervision for next-token prediction for the multimodal LLM.
3.2.2. Multi-task heads
@@
(2) $\mathcal{L}$ $=\mathcal{L}_{\textrm{Next-token}}+\mathcal{L}_{\textrm{Volume}}+\mathcal{L}_{\textrm{Region}}$
$\quad+\mathcal{L}_{\textrm{Shape}}+\mathcal{L}_{\textrm{Spread}}+\mathcal{L}_{\textrm{Out-of-scope}}$
4. Experiments
4.1. Datasets details
@@
For our synthetic VQA protocol, we leverage the 2024 Brain Tumor Segmentation (BraTS) challenge (LaBella et al., 2024), which provides a standardized benchmarking environment for automated brain tumor segmentation. All datasets comprise of co-registered multiparametric MRI scans (T1, T1Gd, T2, FLAIR) at 1mm3 resolution, skull-stripped and manually annotated by experts. To enable fair comparison and manage GPU memory, all BraTS sequences were resampled to $32\times 256\times 256$. This allowed for compatibility with baseline methods, such as M3D (Bai et al., 2024) and Med3DVLM (Xin et al., 2025). We consider three challenges in BraTS: GLI, MET and GoAT. The challenges are collected from over ten institutions and encompass diverse pathological contexts and imaging protocols.
GLI (Adult Glioma Post Treatment)
@@
Dataset focuses on post-treatment diffuse glioma segmentation and consists of multi-institutional routine post-treatment clinically-acquired multiparametric mpMRI scans of glioma. The task requires the delineation of enhancing tumor (ET), non-enhancing tumor core (NETC), surrounding FLAIR hyperintensity (SNFH), and resection cavity (RC) (de Verdier et al., 2024).
MET (Brain Metastases)
@@
Dataset contains a retrospective compilation of treatment-naive brain metastases mpMRI scans obtained from various institutions under standard clinical conditions. The challenge addresses the segmentation of small metastatic lesions using a 3-label system (NETC, SNFH, ET) and demonstrates variable tumor component distribution across cases (Moawad et al., 2024).
GoAT (Generalizability Across Tumors)
@@
Dataset assesses algorithmic generalizability across different tumor types (i.e., different number of lesions per scan, lesion sizes, and locations in the brain), institutions (i.e., different MRI scanners, acquisition protocols), and demographics (i.e., different age, sex, etc.). The challenge uses consistent labels (necrosis, edema/invaded tissue, and enhancing tumor) despite varying tumor morphology to evaluate algorithm adaptability to new disease types with limited training data (de Verdier et al., 2024; Moawad et al., 2024; LaBella et al., 2023; Kazerooni et al., 2024; Adewole et al., 2023).
@@
To generate the train, validation, and test sets, we randomly sample 80%, 10%, and 10% from the imaging studies. For GLI we generated 31,104, 4,176, and 3,624 question-answer pairs for the train, validation, and test sets based on 1,621 mpMRIs. For MET, we generated 9,090, 1,368, and 1,260 question-answer pairs for the train, validation, and test sets, based on 651 mpMRIs. For GoAT, we generated 19,440, 2,430, and 2,448 question-answer pairs for the train, validation, and test sets, based on 1351 mpMRIs. All the generated questions pass a rules-based filter which checks for proper formatting. Additionally, all the generated test set questions pass an automated data cleaning step using ChatGPT-4o to ensure high quality evaluation.
Clinical validation
We collaborated with two radiologists who annotated 20 mpMRIs from the BraTS-GLI test set, 10 mpMRIs from the BraTS-MET test set, and 10 mpMRIs from the BraTS-GoAT test set with questions spanning four tasks and four findings for the BraTS-GLI dataset and three findings for the BraTS-MET and BraTS-GoAT datasets, yielding 560 questions. We used 10 annotated mpMRIs from BraTS-GLI for validation to improve the task label thresholds for the synthetic data. We used the other 30 annotated mpMRIs to evaluate the inter-annotator agreement between the synthetic data and each radiologist, the model predictions and each radiologist, and both radiologists (Table 15 and 16).
To assess the quality of our synthetic questions, a radiologist evaluated the clarity of 160 synthetic questions from 10 mpMRIs from the BraTS-GLI test set using a binary scoring system (1 = valid, 0 = invalid). Prior to the LLM automated data cleaning, the synthetic questions achieved a 92.2% validity rate. After automated data cleaning, the synthetic questions achieved a 95.9% validity rate. More details about the datasets can be found in Appendix A and  B.
4.2. Experimental settings
@@
We compare our approach to several baseline models, including LLaVA-Med (Li et al., 2023a), M3D (Bai et al., 2024), Merlin (Blankemeier et al., 2024), Med3DVLM (Xin et al., 2025), and RadFM (Wu et al., 2023). To process the multiple 3D MRI image modalities, we use a multi-image approach, in which we concatenate the image tokens generated from each MRI image modality from a shared projection layer and vision encoder (Wu et al., 2023). Because LLaVA-Med is not implemented with a 3D vision encoder, to ensure a fair comparison, we test it with our model’s vision encoder (Bai et al., 2024). Similar to our method, the vision encoder is frozen and only the projection layer and LLM are trainable. To provide a comparison to our model’s multi-task classifier, which is trained end-to-end with the rest of our framework, we independently train a new multi-task classifier. We use a Phi3 language model with multi-task heads to predict the multi-task outputs given the prompt and text generation. The model is trained on our train dataset and had 99.8% accuracy on the validation set. Other hyperparameter settings mirror our method as closely as possible to ensure a fair comparison.
@@
All our experiments were mainly conducted using a single NVIDIA A100 GPU on an internal cluster. Training our model on the GLI dataset took approximately 10 hours.
Table 2. Ablation study on the MoE architecture on the GLI validation set with accuracy metric.
@@
rowsep=0.95pt Modality-level MoE Token-level MoE Prompt-based MoE weights Task Mean 63.3 64.1 64.4 65.5
Table 3. Comparison of MoE-based and shared expert approach based on task on the GLI validation set with accuracy metric
Method Volume Region Shape Spread
Shared expert 53.2 82.8 59.7 57.5
MoE-based 57.6 83.5 60.9 59.9
Table 4. Comparison of MoE-based and shared expert approach based on finding on the GLI validation set with accuracy metric
Method ET SNFH NETC RC
Shared expert 63.2 70.8 71.3 48.7
MoE-based 65.1 72.5 75.1 49.9
Table 5. Comparison of MoE approaches based on task on the GLI validation set with accuracy metric
Method Volume Region Shape Spread
Modality-level MoE 53.7 83.4 59.7 59.4
Token-level MoE 56.0 83.5 60.0 58.1
Prompt-based MoE 57.6 83.5 60.9 59.9
Table 6. Comparison of MoE approaches based on finding on the GLI validation set with accuracy metric
Method ET SNFH NETC RC
Modality-level MoE 63.5 71.1 73.9 48.6
Token-level MoE 64.0 72.2 72.2 49.9
Prompt-based MoE 65.1 72.5 75.1 49.9
Table 7. Comparison of task means with and without region information on the GLI validation set with accuracy metric.
Model Task Mean without Region Scores
Without Region Information 59.5
With Region Information 64.1
Table 8. Radiologist acceptance rate comparison between mpLLM and M3D.
Model Radiologist Acceptance Rate (%)
M3D 34.1
mpLLM 50.0
Table 9. Comparison of model performance for differentiating primary gliomas versus secondary metastatic lesions
Model Accuracy AUROC
M3D 88.5 95.5
mpLLM 95.6 99.0
4.3. Results
All model results across the evaluated datasets are presented in Table 1. Our model consistently achieves strong performance across all task categories and datasets, outperforming the second-best model by an average margin of 5.5 points (+9.1% relative to the strongest baseline). We note a small drop in performance relative to other models on the GoAT dataset. This dataset is extremely challenging because it evaluates generalizability based on several factors with limited training data. In our experiments, we also noticed that the top three models had above a 99.8% accuracy on out-of-scope task identification, which suggests our dataset was effective at hallucination mitigation.
4.4. Ablation studies
An ablation study on the MoE architecture is in Table 2. Image modality-level and token-level high-level MoE experts perform better than the single projection layer baseline approach. A prompt-conditioned weighted combination of the different high-level experts performs the best.
Fine-grained results comparing our MoE-based approach and a single shared expert are in Table 3 and Table 4. The more complex multimodal reasoning is helpful for all tasks and findings, and especially helpful for the volume task (+4.4 points and 8.3% relative to the baseline) and NETC finding (+3.8 points and 5.3% relative to the baseline). Fine-grained results comparing modality-level MoE, token-level MoE, as well prompt-conditioned MoE on tasks as well as findings are in Table 5 and Table 6. Token-level MoE is stronger in the volume, region, and shape tasks while modality-level MoE is stronger in the spread task. For findings, token-level MoE is stronger with the ET, SNFH, and RC findings while modality-level MoE is stronger with the NETC finding. The prompt-conditioned MoE excels in all of them, suggesting that based on the question, the model is able to accurately combine the optimal token-level MoE or modality-level MoE blocks.
4.5. Clinical perspective
From a neuroradiology standpoint, an important use case for mpLLM is as an adjunct to clinical expertise: producing standardized, interpretable tumor descriptors (e.g., size, morphology, extent, and coarse localization) for referring neurosurgeons. This framework also provides a pathway to more fine-grained future outputs, such as involvement with specific gyri or eloquent cortex. A unique technical challenge in neuro-oncology imaging is robustness to atypical anatomy and our results suggest the proposed approach remains effective across highly variable lesion appearances.
To quantify clinical utility under this framing, we evaluate (i) agreement between model predictions and radiologist annotations, (ii) radiologist acceptability of generated responses, and (iii) clinically relevant downstream performance. Agreement results are shown in Table 16. We observe that mpLLM is comparable to radiologist-level consistency on the volume and shape tasks, indicating strong reliability for two core descriptors commonly communicated in clinical workflows. Performance is lower on region and spread, which we attribute primarily to higher noise in the synthetic labeling pipeline for these attributes and to the inherent difficulty of localization/extent estimation when anatomy is atypical or deformed. This result highlights a practical constraint for real-world deployment: model performance is bounded by label quality and by the stability of the underlying definitions.
We further probe the role of localization by appending region information to the prompt (Table 7), yielding an approximately 5% increase on other task scores. This suggests (i) localization provides useful signal for estimating morphology/extent and (ii) even relatively noisy region labels can improve other clinically relevant descriptors, warranting their inclusion. This aligns with clinical practice, where coarse localization is often communicated alongside morphology/extent.
Because many descriptors are subjective, we collaborated with a radiologist to assess whether generated answers are clinically reasonable. We created a clinical validation set of 208 questions from 13 GLI test cases, focusing on volume, region, shape, and spread (questions 1–4 in Appendix A). We compared mpLLM against M3D by asking the radiologist to label each response as sufficient or lacking. mpLLM achieves a 50.0% acceptability rating and improves clinical acceptability by +15.9 points (+46.6% relative; Table 8). These findings support the role of mpLLM as a reporting adjunct for tumor descriptors.
Finally, to emphasize downstream usefulness, we consider a clinically important task: differentiating primary gliomas from secondary metastatic lesions using imaging-derived descriptors (volume, region, shape, spread). Because our models are trained with guard rails and indicate if tasks are outside of their scope, we fit a logistic regression classifier on model-generated features from the combined training set of BraTS GLI and MET. On the test set (Table 9), mpLLM improves over M3D by +7 accuracy points and +4 AUROC points, suggesting that the learned structured descriptors capture discriminative imaging patterns relevant to neuro-oncology workflows.
5. Conclusion
We present mpLLM, a multimodal LLM for mpMRI VQA that uses prompt-conditioned hierarchical MoE routing over modality- and token-level projection experts, enabling efficient end-to-end fine-tuning without paired image–report pretraining. We pair this model with a synthetic VQA pipeline derived from expert segmentation annotations, developed in collaboration with a radiologist, to produce standardized tumor descriptors (e.g., volume, morphology, extent, and coarse localization) that can serve as an adjunct to clinical expertise for referring neurosurgeons. Across BraTS GLI/MET/GoAT, mpLLM improves over strong medical VLM baselines by an average of +5.5 points and increases radiologist-rated clinical acceptability by +15.9 points, with ablations validating the contributions of modality/token experts and prompt-conditioned routing. These findings suggest that structured, interpretable outputs can remain reliable across heterogeneous lesion appearances, while also highlighting the importance of label quality for more challenging localization tasks. Future work includes extending to open-ended VQA and report generation, evaluating robustness in settings with atypical anatomy, conducting broader multi-reader validation, and performing fairness analyses.
6. Limitations and Ethical Considerations
This work uses publicly available, fully de-identified BraTS datasets, minimizing risks to patient privacy and data security. Our synthetic VQA are generated from segmentation annotations, and both the generated questions and model outputs underwent radiologist review to mitigate typical risks of synthetic supervision. Nonetheless, fairness and bias remain open concerns: synthetic prompts and limited demographic metadata can yield models that underperform for underrepresented groups or clinical scenarios. The model is intended for research only and must not be used for autonomous clinical decision-making; it is designed to abstain on out-of-scope queries, and any deployment would require prospective, multi-site validation under qualified clinical oversight. In future work, we will expand evaluations to demographically diverse cohorts where available, document dataset composition and known limitations, and incorporate explicit fairness analyses and bias-mitigation strategies alongside robustness and calibration assessments.
7. GenAI Disclosure
Appendix A Additional information regarding dataset
@@
Table 10. Statistics for the synthetic VQA datasets.
@@
In the following, we will describe the formulas used to derive the shape and spread descriptors for our synthetic VQA protocol. Let $M\subset\mathbb{Z}^{3}$ be a binary mask of lesion voxels sampled with spacing $\mathbf{s}=(s_{x},s_{y},s_{z})\,[\mathrm{mm}]$ (typically $s_{x}=s_{y}=s_{z}=1$). Write $\Delta V=s_{x}s_{y}s_{z}$ for the physical volume of one voxel and $\lvert M\rvert$ for the number of lesion voxels.
@@
We define 26-connectivity on $M$ (scipy ‘ndimage.label‘ with a unit “ball” structuring element) and record $N_{\!c}$ and $M_{1},\dots,M_{N_{\!c}}$.
@@
$\text{shape}=\begin{cases}\text{``focus''}&V_{\text{tot}}<0.1\;\text{cm}^{3}\quad(\;V_{\text{tot}}\times 10^{-3}<0.1\;)\\[2.0pt] \text{``round''}&\Phi\geq 0.80\;\land\;E\leq 1.4,\\[2.0pt] \text{``oval''}&0.50\leq\Phi<0.80\;\land\;1.4<E\leq 3.0,\\[2.0pt] \text{``elongated''}&E>3.0,\\[2.0pt] \text{``irregular''}&\text{otherwise.}\end{cases}$
The thresholds were set empirically on 10 annotated mpMRIs from BraTS-GLI for validation set and match radiologists’ qualitative intuition of near-spherical, mildly flattened, and strongly stretched geometries. All computations are implemented in Python using scipy, scikit-image, numpy, and ndimage as shown in the listing above.
@@
We use ChatGPT to generate question augmentations of our multitask dataset. For generating question augmentations for the standard multi-task prompts, we first provide this prompt “Please produce hundred alternative wordings that a radiologist may use for the following question and answer. Please include everything surrounded by curly braces {} as they are because they are placeholders. Please generate the reworded question starting with “Q:” and reworded answer starting with “A:” and separate each generated question-answer pair with a newline. Please do not produce any additional text.” and append this to each of the multitask prompts below. We produce 40 repetitions with a temperature of $1.0$, top p of $1$, and model “gpt-4o-mini-2024-07-18”.
  1. (1)
@@
For generating question augmentations for the partially out-of-scope multi-task prompts, we first provide this prompt “Please produce hundred alternative wordings that a radiologist may use for the following question and answer and incorporate an additional clinical task or tasks which the model cannot solve in the reworded question. These can be before, after, or interspersed between the other tasks (please make sure to vary the order and number of out-of-scope tasks). Do not mention that the model cannot answer these in the question; however, indicate that the model cannot answer that part of the question in the reworded answer (potentially using different phrasings). The model can describe the volume, brain region, shape, and spread of {label} which is the region of interest. Please include everything surrounded by curly braces {} as they are because they are placeholders. Please generate the reworded question starting with “Q:” and reworded answer starting with “A:” and do not produce any additional text.” and append this to each of the multitask prompts above. We produce 10 repetitions with a temperature of $1.0$, top p of $1$, and model “gpt-4o-2024-08-06”.
@@
For generating question augmentations for completely out-of-scope prompts, we first provide this prompt “Please produce a hundred questions (with one or more tasks) that a radiologist may ask that the model does not have information to answer. The model can describe the volume, brain region, shape, and spread of {label} which is the region of interest. Please include {label} in the question but do not include anything else with curly braces. In the answer, please indicate the model cannot answer the question (potentially using different phrasings). Please generate the question starting with “Q:” and answer starting with “A:” and do not produce any additional text.” We produce 10 repetitions with a temperature of $1.0$, top p of $1$, and model “gpt-4o-mini-2024-07-18”.